OpenStudy (bruised1998):

5-3^x=-40 round to the nearest ten-thousandth (gradpoint)

OpenStudy (3mar):

If you want to solve for x, we would proceed like that. \[5-3^x=-40\] \[5-3^x+40=-40+40\] \[45-3^x=0\] \[45-3^x+3^x=0+3^x\] \[45=3^x\] now you can take the log for both sides \[\log(45)=\log(3^x)\] \[\log(45)=xlog(3)\] \[\frac{ \log(45) }{ \log(3) }=x\] \[x=\frac{ \log45 }{ \log3 }=3.465\] I hope I helped you.

OpenStudy (bruised1998):

thank you for showing the problem out like this

OpenStudy (3mar):

You are welcome! Pleasure is mine!